OpenStudy (anonymous):

find the indefinite integral: ∫3x^(x)sin(x^(3))dx

OpenStudy (anonymous):

\[\int\limits3x ^{2}\sin x ^{3}dx\]

OpenStudy (zarkon):

\[u=x^3\]

OpenStudy (anonymous):

i know...but i want to check my ans...\[\frac{ -\cos x^ 4 }{ 4 }\] ?

OpenStudy (shamim):

i think answer will b \[-\cos x ^{3}\]

OpenStudy (shamim):

can i show details

OpenStudy (jhannybean):

Did you use integration by parts?

OpenStudy (shamim):

i can show my work

OpenStudy (anonymous):

oops ...can you show me the detail...? thanks

OpenStudy (shamim):

ok

OpenStudy (shamim):

let \[u=x ^{3}\]

OpenStudy (shamim):

or,\[du=3x ^{2}dx\]

OpenStudy (shamim):

so ur given term will b\[\int\limits_{?}^{?} \sin u du=-\cos u=-cosx ^{3}\]

OpenStudy (jhannybean):

You can take out the 3 since its just a constant multiplied to the whole thing, so your \[u=x^3\]

OpenStudy (shamim):

no need to take 3 outside of ur integral

OpenStudy (jhannybean):

-_- i read it as 3x^3 xD my bad.I got the same answer.:)